# Why Smishing Works

### v2

@@ 4.4.2. Second Round Questions
In the second round, participants were presented with their responses to each ”how likely they were to click on one of the links, reply to the text, or call the number” first round question, along with the respective screenshot. Participants were asked again, ”How likely are you to interact with this SMS?” with the same options as in the first round. This was used to measure if after seeing the message a second time caused them to change their mind. Finally, we asked participants, ”Do you think this SMS is real or fake?” This process repeated for all 16 messages. A similar approach has been used in prior phishing related user study tasks (neupane2015multi).
@@ 4.5. Post-Survey
Participants completed five post-survey questionnaires to comprehensively assess the participants’ security and attention behavior. Initially, participants responded to the RSeBIS questionnaire, a revised version of the Security Behavior Intentions Scale (SeBIS) (sawaya2017self; egelman2015scaling). This questionnaire gauges user security approaches, encompassing topics like password creation and device updates. Additionally, we aimed to understand participant attitudes toward utilizing security tools, administering the Security Attitudes (SA-6) questionnaire. Comprising six questions, the SA-6 is designed to measure user security behavior and willingness to adopt security methods (faklaris2019self).
Participants also completed an Internet Users’ Information Privacy Concerns (IUIPC) questionnaire to gain insight into their security behavior and perceptions of online threats. To provide a more holistic view of individuals and their relation to cybersecurity, we included a questionnaire on common security terms, such as malware and phishing. Participants indicated their familiarity with these terms on a five-point scale ranging from ‘not familiar at all’ to ‘extremely familiar.’ Recognizing that understanding these terms requires cognitive capacity, we included the Attention Control Scale (ATTC) questionnaire to measure user attention control when interacting with security concepts (derryberry2002anxiety).
@@ 4.6. Data Quality and Attention Checks
In the initial survey collection, we received 423 responses. To uphold data quality, we implemented various checks. Initially, we filtered responses to only include participants with a US GeoIP location, restricting it to one completion per participant IP. Subsequently, we excluded participants who did not fully complete the survey. As an additional measure to enhance data quality, we scrutinized respondents with completion times significantly lower than the average. To identify these outliers, we utilized the median absolute deviation (MAD) of the overall survey response time, opting for the median over the mean to mitigate the impact of variance at edge cases. Establishing a threshold of one MAD from the median response time, we marked responses below this threshold as outliers, subsequently eliminating them from our study. Following these quality checks, we retained 214 participants for our analysis. We performed a power analysis for
Throughout our survey procedure, we incorporated five attention check questions to gauge participant attention and question comprehension. These questions encompassed two widely used forms: instructed response and instructional manipulation check (IMC) (kung2018attention). Correct answers to these questions were tallied, and each participant received a score from one to five. We then compared participants’ relative attention scores to their accuracy for real, fake, and overall messages. The average accuracy rates at each score level are presented in Table LABEL:tab:att-score-accuracy in the appendix. To assess differences in accuracy between attention scores, a one-way ANOVA was conducted. No significant difference was found in mean overall accuracy scores when comparing participants’ attention totals (F(54, 143) = 0.965, p = 0.549). Consequently, we retained all remaining responses, irrespective of their attention check score.
@@ 5.1. Answering RQ:1
Initially, we analyzed the accuracy rates of the messages, as illustrated in Figure LABEL:tab:accuracy-rates-bymessage in the appendix. This table presents the average rates at which participants correctly identified our messages as real or fake. The accuracy rates per message, as well as the overall accuracy for both real and fake messages, are depicted.
For fake SMS, the average accuracy per message ranged from 60.2% to 71.5%, with an overall average accuracy of 65.6%. In terms of real SMS, our accuracy ranged from 34.3% to 53.2%, with an average accuracy of 44.6%.
Next, we assessed how a user’s ability to accurately distinguish between real and fake messages was influenced by the metrics derived from the 16 questions in the smishing detection round. The average rates for each response across the first and second rounds are depicted in Figure 1.
@@
Participants were asked twice regarding their likelihood to interact with the messages. A comparison of interaction likelihood between rounds one and two revealed a decrease of 0.11 points for fake messages and 0.10 points for real messages on the six-point Likert scale. Nevertheless, when examining the mean scores of these 214 participants, no significant difference was found in the likelihood to interact with fake messages (t(424) = 0.900, p = 0.369) or real messages (t(423) = 0.881, p = 0.379).
Figure 2. Comparison of correct answers against brand familiarity, interaction rates and whether they have received a message from the brand before.
@@
In Figure 1, we also present the frequency of each answer for the first round of questions with the SMS stimuli and our second round questions, categorized by real and fake messages. The mean Likert scores for the likelihood of interacting with real messages were 2.90, while the scores for fake messages were 2.77. Overall, this indicates that participants were slightly more inclined to interact with real messages than fake ones. The same trend holds true for comfortability with interacting with the messages, with a 2.71 average Likert score for real messages and 2.63 for fake messages. Participants scored brand familiarity with an average of 3.00 for fake messages and 3.02 for real messages, indicating a similar level of familiarity with both. Additionally, the percentage of participants who had received an SMS from the brand before was similar between real and fake messages. Lastly, we present the real and fake responses to the smishing detection task, revealing numerous false positive and false negative results. Notably, among the last two questions, a significant percentage of ‘I don’t know’ responses indicate uncertainty about whether participants had received a message before and whether the messages were real or fake.
@@
Subsequently, we conducted a linear regression analysis on the interaction metrics described in Figure 1 to determine whether the metrics from our first-round smishing detection task significantly predicted participants’ accuracy in identifying the message as real or fake. For each case, we tested the null hypothesis with the F-test for overall regression.
@@
The linear regression for fake messages was statistically significant ($R^{2}$ = 0.477, F(4, 189) = 43.10, p ¡ 0.001). The regression analysis revealed that the likelihood of interacting with fake messages significantly predicted accuracy with fake messages ($\beta$ = -0.52, p ¡ 0.001). However, familiarity with the brand, previous receipt of an SMS from the brand, and comfortability with interacting did not significantly predict smishing detection accuracy for fake messages.
@@
Similarly, for real messages, the linear regression was statistically significant ($R^{2}$ = 0.376, F(4, 189) = 28.14, p ¡ 0.001). The regression analysis demonstrated that the likelihood of interacting with real messages significantly predicted accuracy with real messages ($\beta$ = 0.294, p = 0.008). Additionally, comfortability with interacting also significantly predicted accuracy with real messages ($\beta$ = 0.323, p = 0.004). However, familiarity with the brand and previous receipt of an SMS from the brand did not significantly predict smishing detection accuracy for real messages. To illustrate the trend of likelihood rates, familiarity, and receipt as accuracy increases, we provide a plot in Figure 2 showing the average scores at each interval of the number of correct answers to determine whether the message is real or fake.
@@ IV F p $\eta^{2}$ 95% CI F p $\eta^{2}$ 95% CI
Sex 1.86 .158 0.019 [0.00-0.07] 0.55 .575 0.006 [0.00-0.04]
Ethnicity 2.28 .038 0.067 [0.00-0.12] 1.55 .163 0.047 [0.00-0.09]
Age 2.41 .051 0.048 [0.00-0.10] 1.74 .144 0.035 [0.00-0.08]
Income 2.19 .037 0.075 [0.00-0.12] 1.49 .174 0.053 [0.00-0.09]
Smartphone Use 1.56 .186 0.032 [0.00-0.08] 0.43 .784 0.009 [0.00-0.03]
Education 0.74 .653 0.031 [0.00-0.05] 0.94 .479 0.034 [0.00-0.06]
Occupation 2.14 .019 0.114 [0.02-0.15] 1.40 .178 0.077 [0.00-0.10]
@@
Sex 2.19 .115 0.020 [0.00-0.07] 1.13 .326 0.011 [0.00-0.05]
Ethnicity 3.43 .003 0.090 [0.01-0.15] 2.78 .013 0.075 [0.00-0.13]
Age 5.10 <.001 0.089 [0.02-0.15] 2.72 .031 0.050 [0.00-0.10]
Income 4.98 <.001 0.145 [0.04-0.21] 3.79 <.001 0.115 [0.02-0.17]
Smartphone Use 0.37 .824 0.007 [0.00-0.02] 1.26 .287 0.024 [0.00-0.06]
Education 0.61 .771 0.023 [0.00-0.03] 1.10 .368 0.041 [0.00-0.07]
Occupation 2.58 .004 0.120 [0.01-0.17] 0.91 .529 0.048 [0.00-0.06]
@@
Sex 2.09 .126 0.020 [0.00-0.06] 0.49 .613 0.005 [0.00-0.03]
Ethnicity 5.79 <.001 0.144 [0.05-0.21] 4.01 <.001 0.105 [0.02-0.16]
Age 4.14 <.001 0.092 [0.02-0.16] 2.81 .026 0.051 [0.00-0.10]
Income 4.74 <.001 0.139 [0.04-0.20] 5.05 <.001 0.147 [0.05-0.21]
Smartphone Use 0.93 .446 0.018 [0.00-0.05] 0.90 .462 0.017 [0.00-0.05]
Education 0.64 .746 0.024 [0.00-0.04] 0.64 .740 0.025 [0.00-0.04]
Occupation 2.61 .004 0.125 [0.02-0.17] 0.88 .564 0.046 [0.00-0.06]
@@ 5.2. Answering RQ:2
To investigate the roles of demographics and behavioral predictors in influencing users’ susceptibility to smishing attacks, we conducted an ANOVA test. This test aimed to compare the impacts of various demographic factors on the likelihood of users to interact with both real and fake messages, as well as their accuracy in identifying messages. The ANOVA test results, along with their effect sizes, are presented in Table 1. For a detailed breakdown of our demographics, please refer to Table LABEL:tab:DemographicAttribute2 in the appendix. Given that we do not assume homogeneity of variances in our responses, we employed a Games-Howell post hoc test for multiple comparisons when the results of our ANOVA test were deemed significant.
@@ 5.2.1. Ethnicity
Comfortability with interacting: The post hoc test results show that the average comfort level with interacting with fake messages was significantly lower for Hispanic participants in comparison with White (p $<$ 0.001, 95% C.I. = [-1.75, -0.91]), Asian (p = 0.019, 95% C.I. = [-2.05, -0.14]), Black (p $<$ 0.001, 95% C.I. = [-2.17, -0.51]) and Native American (p $<$ 0.001, 95% C.I. = [-2.03, -0.84]). The post hoc test results also show that the average comfort level with interacting with real messages was significantly lower for Hispanic participants in comparison with White (p = 0.039, 95% C.I. [-1.90, -0.04]), Black (p = 0.013, 95% C.I. = [-2.34, -0.20]) and Native American (p = 0.038, 95% C.I. = [-1.98, -0.04]).
Accuracy: The post hoc test results show the average accuracy with fake messages was significantly higher for Hispanic participants in comparison with White (p $<$ 0.001, 95% C.I. = [0.16, 0.43]), Black (p = 0.007, 95% C.I. = [0.10, 0.78]) and Native American (p = 0.015, 95% C.I. = [0.06, 0.70]).
@@ 5.2.2. Age
Likely to interact: The post hoc test results show that the average likely interaction rates with fake messages was significantly lower for participants in the 18-24 age category were statistically different than those in the 25-34 (p = 0.034, 95% C.I. = [-1.36, -0.04]) and 35-44 (p $<$ 0.001, 95% C.I. = [-1.87, -0.40]) category. The post hoc test results show that average likely interaction rates with real messages were not significant.
Comfortability with interacting: The post hoc test results show that the average comfort levels with fake messages was significantly lower for participants in the 18-24 age category were statistically different than those in the 35-44 (p = 0.002, 95% C.I. = [-1.49, -0.26]) and 45-54 (p = 0.004, 95% C.I. = [-1.50, -0.23]) age category. The post hoc test results also show that the average comfort levels with real messages was significantly higher in participants in the 45-54 age category than those in the 18-24 (p = 0.026, 95% C.I. = [0.05, 1.15]) and 25-34 (p $<$ 0.001, 95% C.I. = [0.20, 0.62]) age category.
@@ 5.2.3. Income level
Likely to interact: The post hoc test results show that the average likely interaction rates with fake messages was significantly lower for participants with an income less than \$10,000 than participants in the \$20,000-\$39,999 (p = 0.028, 95% C.I. = [-3.12, -0.15]), \$40,000-\$59,999 (p = 0.020, 95% C.I. = [-3.21, -0.28]) and \$40,000-\$59,999 (p = 0.032, 95% C.I. = [-3.05, -0.12]) range. The post hoc results for average likely interaction rates with real messages were not significant.
Comfortability with interacting: The post hot test results show that the average comfort levels with fake messages was significantly lower for participants with an income less than \$10,000 than participants in the \$20,000-\$39,999 (p = 0.014, 95% C.I. = [-1.80, -0.16]), \$40,000-\$59,999 (p = 0.006, 95% C.I. = [-1.85, -0.31]) and \$40,000-\$59,999 (p = 0.010, 95% C.I. = [-1.77.-0.21]) range. The post hoc results for average likely interaction rates with real messages were not significant.
Accuracy: The post hoc test results show that the average accuracy rates fake messages was significantly higher for participants with an income less than \$10,000 than participants in the \$20,000-\$39,999 (p = 0.041, 95% C.I. = [0.01, 0.60]), \$40,000-\$59,999 (p $<$ 0.001, 95% C.I. = [0.14, 0.59]) and \$40,000-\$59,999 (p $<$ 0.001, 95% C.I. = [0.14, 0.64]) range.
5.2.4. Sex, Smartphone Use, Education Levels

### v4

@@ 4.4.2. Second Round Questions
In the second round, participants were then presented with their responses to each ”how likely they were to click on one of the links, reply to the text, or call the number” first round question, along with the respective screenshot. Participants were asked again, ”How likely are you to interact with this SMS?” with the same options as in the first round. This was used to measure if after seeing the message a second time caused them to change their mind. Finally, we asked participants, ”Do you think this SMS is real or fake?” This process repeated for all 16 messages. A similar approach has been used in prior phishing related user study tasks (neupane2015multi, ).
@@ 4.5. Post-Survey
Participants completed five post-survey questionnaires to comprehensively assess the participants’ security and attention behavior. Initially, participants responded to the RSeBIS questionnaire, a revised version of the Security Behavior Intentions Scale (SeBIS) (sawaya2017self, ; egelman2015scaling, ). This questionnaire gauges user security approaches, encompassing topics like password creation and device updates. Additionally, we aimed to understand participant attitudes toward utilizing security tools, administering the Security Attitudes (SA-6) questionnaire. Comprising six questions, the SA-6 is designed to measure user security behavior and willingness to adopt security methods (faklaris2019self, ).
Participants also completed an Internet Users’ Information Privacy Concerns (IUIPC) questionnaire to gain insight into their security behavior and perceptions of online threats (Gross2021ValidityAR, ). To provide a more holistic view of individuals and their relation to cybersecurity, we included a questionnaire on common security terms, such as malware and phishing. This survey utilizes a similar approach to related works (10.5555/3563609.3563635, ), and was expanded to include common terms relating to SMS security from the NCSC glossary (NSCS2022, ). Participants indicated their familiarity with these terms on a five-point scale ranging from ‘not familiar at all’ to ‘extremely familiar.’ Recognizing that understanding these terms requires cognitive capacity, we included the Attention Control Scale (ATTC) questionnaire to measure user attention control when interacting with security concepts (derryberry2002anxiety, ). After the post survey, participants were debriefed about the nature of the study.
@@ 4.6. Data Quality and Attention Checks
In the initial survey collection, we received 423 responses. To uphold data quality, we implemented various quality checks in our survey. We used the Qualtrics filtering option to limit responses to only include participants with a US GeoIP location and enabled the prevent multiple submissions option. Throughout the survey we also incorporated five attention check questions to gauge the participant attention and question comprehension. These questions encompassed two widely used forms: instructed response and instructional manipulation check (IMC) (kung2018attention, ; commitqualtrics, ). Subsequently, we excluded participants who did not fully complete the survey or who failed more than two of the attention check questions we spread throughout the questionnaires. Following these quality checks, we retained 187 participants for our analysis.
@@ 5.1. Answering RQ:1
Initially, we analyzed the accuracy rates of the messages, as illustrated in Figure 3 in the appendix. This table presents the average rates at which participants correctly identified our messages as real or fake. The accuracy rates per message, as well as the overall accuracy for both real and fake messages, are depicted. The average accuracy per fake message ranged from 63.1% to 73.5%, with an overall average accuracy of 67.1%. For real messages, our accuracy ranged from 34.0% to 51.8%, with an average of 43.6%.
@@
Next, we assessed how a user’s ability to accurately distinguish between real and fake messages was influenced by metrics related to the 16 questions in the smishing detection round. The average rates for each response across the first and second rounds are depicted in Figure 1. Participants were asked twice regarding their likelihood to interact with the messages. A comparison of interaction likelihood between rounds one and two revealed a decrease of 0.12 points for fake messages and 0.11 points for real messages on the six-point Likert scale. Nevertheless, when examining the mean scores of these 187 participants, we found no significant difference in the likelihood to interact with fake messages (t(370) = 0.64, p = 0.425) or real messages (t(369) = 1.54, p = 0.216).
@@
In Figure 1, we also present the frequency of each answer for the first round of questions with the SMS stimuli and our second round questions, categorized by real and fake messages. The mean Likert scores for the likelihood of interacting with real messages were 2.83, while the scores for fake messages were 2.70. Overall, this indicates that participants were slightly more inclined to interact with real messages than fake ones. The same trend holds true for comfortability with interacting with the messages, with a 2.67 average Likert score for real messages and 2.47 for fake messages. Participants scored brand familiarity with an average of 3.02 for fake and real messages, indicating a similar level of familiarity with both. Similarly, participants reported having received messages from the brands listed in the messages an average 55% of the time for both real and fake messages. Lastly, we present the real and fake responses to the smishing detection task, revealing numerous false positive and false negative results. Notably, among the last two questions, a significant percentage of ‘I don’t know’ responses indicate uncertainty about whether participants had received a message before and whether the messages were real or fake.
@@
Subsequently, we conducted a linear regression analysis on the interaction metrics described in Figure 1 to determine whether the metrics from our first-round smishing detection task significantly predicted participants’ accuracy in identifying the message as real or fake. For each case, we tested the null hypothesis with the F-test for overall significance.
@@
The linear regression for fake messages was statistically significant ($R^{2}$ = 0.518, F(4, 166) = 46.73, p ¡ 0.001). The regression analysis revealed that the likelihood of interacting with fake messages significantly predicted accuracy with fake messages ($\beta$ = -0.679, p ¡ 0.001). However, familiarity with the brand, previous receipt of an SMS from the brand, and comfortability with interacting did not significantly predict smishing detection accuracy for fake messages.
@@
Similarly, for real messages, the linear regression was statistically significant ($R^{2}$ = 0.374, F(4, 164) = 28.12, p ¡ 0.001). The regression analysis demonstrated that the likelihood of interacting with real messages significantly predicted accuracy with real messages ($\beta$ = 0.387, p = 0.002). Additionally, comfortability with interacting also significantly predicted accuracy with real messages ($\beta$ = 0.251, p = 0.037). However, familiarity with the brand and previous receipt of an SMS from the brand did not significantly predict smishing detection accuracy for real messages. Alternatively, we did not find any significant differences in regards to messaging history or familiarity with the brand.
RQ1 Summary: These results highlight the correlation between interaction rates, comfort with interaction, and accuracy rates. Specifically, participants less inclined to engage with fake messages were more adept at correctly identifying them as such. Furthermore, interaction rates and comfort levels were significant predictors of accurately identifying real messages. In contrast, brand familiarity and a history of receiving messages from the brand did not significantly influence accuracy for either real or fake messages. These findings imply that while comfort with interaction may affect the likelihood of engagement, it does not impact the accuracy of identifying fake messages as much as participants’ self-reported likelihood of interaction does.
@@ IV F p $\eta^{2}$ 95% CI F p $\eta^{2}$ 95% CI
Sex 3.39 .036 .038 [.00-.10] 0.97 .382 .011 [.00-.05]
Ethnicity 2.38 .031 .079 [.00-.14] 1.81 .099 .062 [.00-.11]
Age 2.32 .059 .052 [.00-.11] 1.98 .099 .046 [.00-.10]
Income 2.23 .034 .086 [.00-.14] 1.53 .159 .062 [.00-.11]
Smartphone Use 0.17 .955 .004 [.00-.01] 0.41 .800 .010 [.00-.03]
Education .638 .745 .030 [.00-.04] 1.10 .362 .045 [.00-.08]
Occupation 1.57 .113 .097 [.00-.13] 1.26 .253 .080 [.00-.11]
@@
Sex 3.62 .029 .038 [.00-.10] 1.70 .186 .018 [.00-.07]
Ethnicity 4.60 <.001 .133 [.03-.20] 3.08 .007 .094 [.01-.15]
Age 5.66 <.001 .111 [.03-.18] 3.98 .004 .081 [.01-.15]
Income 5.60 <.001 .180 [.06-.25] 3.62 .001 .125 [.02-.19]
Smartphone Use 0.57 .687 .012 [.00-.04] 0.51 .726 .011 [.00-.03]
Education 1.13 .345 .048 [.00-.08] 1.60 .127 .068 [.00-.11]
Occupation 2.17 .018 .120 [.00-.16] 0.68 .752 .041 [.00-.05]
@@
Sex 3.63 .029 .038 [.00-.10] 0.65 .525 .007 [.00-.04]
Ethnicity 6.38 <.001 .176 [.06-.25] 4.18 <.001 .123 [.03-.19]
Age 5.43 <.001 .107 [.02-.18] 3.34 .012 .069 [.00-.13]
Income 5.57 <.001 .180 [.06-.25] 4.80 <.001 .159 [.05-.23]
Smartphone Use 0.25 .911 .005 [.00-.02] 0.48 .750 .011 [.00-.03]
Education 0.81 .597 .035 [.00-.06] 0.99 .444 .043 [.00-.07]
Occupation 2.35 .010 .129 [.01-.17] 0.72 .723 .043 [.00-.05]
@@ 5.2. Answering RQ:2
To investigate the effect of demographics on the behavioral predictors that influence users’ susceptibility to smishing attacks, we conducted an ANOVA test. This test aimed to compare the impacts of various demographic factors on the likelihood of users to interact with both real and fake messages, as well as their accuracy in identifying messages. The ANOVA test results, along with their effect sizes, are presented in Table 1. For a detailed breakdown of our demographics, please refer to Table 5 in the appendix. Given that we do not assume homogeneity of variances in our responses, we employed a Games-Howell post hoc test for multiple comparisons when the results of our ANOVA test were deemed significant. The significance of the results of this test are to inform the understanding of the users’ vulnerabilities to smishing attacks in different populations. Furthermore, by identifying their significant differences in susceptibility to smishing attacks, these findings can allow for targeted future research and resource allocation, directing efforts towards populations most vulnerable.
5.2.1. Sex
Likely to interact: The post hoc test results for average accuracy when identifying fake messages were not significant.
Comfortability with interacting: The post hoc test results show that the average comfort levels with fake messages was significantly lower for female participants in comparison with male participants(p = 0.030, 95% C.I. = [-0.66, -0.03]).
Accuracy: The post hoc test results show that the average accuracy with fake messages was significantly higher for female participants in comparison with male participants(p = 0.032, 95% C.I. = [0.01, 0.26]).
@@ 5.2.2. Ethnicity
Comfortability with interacting: The post hoc test results show that the average comfort level with interacting with fake messages was significantly lower for Hispanic participants in comparison with White (p $<$ 0.001, 95% C.I. = [-1.75, -0.91]), Asian (p = 0.019, 95% C.I. = [-2.05, -0.14]), Black (p = 0.003, 95% C.I. = [-2.24, -0.39]) and Native American (p $<$ 0.001, 95% C.I. = [-2.06, -0.83]). The post hoc test results also show that the average comfort level with interacting with real messages was significantly lower for Hispanic participants in comparison with White (p = 0.033, 95% C.I. [-2.04, -0.07]) and Native American (p = 0.017, 95% C.I. = [-2.41, -0.17]).
Accuracy: The post hoc test results show the average accuracy with fake messages was significantly higher for Hispanic participants in comparison with White (p $<$ 0.001, 95% C.I. = [0.15, 0.43]), Black (p = 0.034, 95% C.I. = [0.02, 0.76]) and Native American (p = 0.011, 95% C.I. = [0.07, 0.73]) participants.
@@ 5.2.3. Age
Likely to interact: The post hoc test results show that the average likely interaction rates with fake messages was significantly lower for participants in the 18-24 age category than those in the 25-34 (p = 0.034, 95% C.I. = [-1.36, -0.04]) and 35-44 (p $<$ 0.001, 95% C.I. = [-1.87, -0.40]) category. The post hoc test results show that average likely interaction rates with real messages were not significant.
Comfortability with interacting: The post hoc test results show that the average comfort levels with fake messages was significantly lower for participants in the 18-24 age category than those in the 35-44 (p = 0.002, 95% C.I. = [-1.49, -0.26]) and 45-54 (p = 0.004, 95% C.I. = [-1.50, -0.23]) age category. The post hoc test results also show that the average comfort levels with real messages was significantly higher in participants in the 45-54 age category than those in the 18-24 (p = 0.026, 95% C.I. = [0.05, 1.15]) and 25-34 (p $<$ 0.001, 95% C.I. = [0.20, 0.62]) age category.
@@ 5.2.4. Income level
Likely to interact: The post hoc test results show that the average likely interaction rates with fake messages was significantly lower for participants with an income less than \$10,000 than participants in the \$20,000-\$39,999 (p $<$ 0.001, 95% C.I. = [-3.08, -1.02]), \$40,000-\$59,999 (p $<$ 0.001, 95% C.I. = [-2.69, -1.32]) and \$80,000-\$99,999 (p $<$ 0.001, 95% C.I. = [-2.11, -0.71]) range. The post hoc results for average likely interaction rates with real messages were not significant.
Comfortability with interacting: The post hoc test results show that the average comfort levels with fake messages was significantly lower for participants with an income less than \$10,000 than participants in the \$20,000-\$39,999 (p = 0.038, 95% C.I. = [-2.08, -0.05]), \$40,000-\$59,999 (p = 0.027, 95% C.I. = [-2.04, -0.13]) and \$80,000-\$99,999 (p = 0.039, 95% C.I. = [-2.00.-0.09]) range. The post hoc results for average likely interaction rates with real messages were not significant.
Accuracy: The post hoc test results show that the average accuracy rates fake messages was significantly higher for participants with an income less than \$10,000 than participants in the \$20,000-\$39,999 (p = 0.023, 95% C.I. = [0.04, 0.72]), \$40,000-\$59,999 (p $<$ 0.001, 95% C.I. = [0.14, 0.57]) and \$80,000-\$99,999 (p $<$ 0.001, 95% C.I. = [0.19, 0.66]) range.
5.2.5. Smartphone Use, Education Levels